Mrs Boyd has a desk full of quarters and dimes. If she has a total of 54 coins with a face value of $12.30. How many of each coin does she have?
Answer by Cromlix(4381)

x = quarters y = dimes
x + y = 54...(1)
25x + 10y = 1230 ....(2)
Multiply (1) by 10
10x + 10y = 540 ....(3)
25x + 10y = 1230....(2)
Subtract (2) from (3)
15x = 690
x = 46
substitute x = 46 into Eq (1)
x + y = 54
46 + y = 54
y = 54 - 46
y = 8
Therefore there are 46 quarters + 8 dimes
Hope this helps
:-)